(i) Show that if A ⊆ C and B ⊆ D, then A×B ⊆ C ×D
(ii) Let A ={a , b , c ,d}and B ={y , z}. Find (a) A×B (b) B×A
1
Expert's answer
2020-07-12T17:29:01-0400

(i) Take "(x,y)\in A\times B". Then "x\in A" and "y\in B", hence by "A\subset C" and "B\subset D" we have "x\in C" and "y\in D." This implies that "(x,y)\in C\times D", which gives "A\times B\subset C\times D".

(ii) By definition, "A\times B=\{(a,y),(a,z),(b,y),(b,z),(c,y),(c,z),(d,y),(d,z)\}" and "B\times A=\{(y,a),(z,a),(y,b),(z,b),(y,c),(z,c),(y,d),(z,d)\}"